OpenStudy (anonymous):

solve -q-11 = 2q + 4

OpenStudy (anonymous):

Start by combining like terms. Move (mathematically ofcourse) all the q's to one side and everything else to the other.

OpenStudy (anonymous):

so if i move it all to one side it is -1q + 2q = -11+4 ?

OpenStudy (anonymous):

is it 7 ?

OpenStudy (anonymous):

-q-11=2q+4 +11 +11 -q=2q+11+4 -2q -2q -q-2q=15 -3q=15

OpenStudy (anonymous):

You can't just move it over, you have to cancel it from the side it was on. Hopefully my work through shows the steps clearly

OpenStudy (anonymous):

oh ok i get it now thanks alot :)